# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ Appendix A Connection to the covariance change detection problems
Wang et al., 2021a proposed a projection-based binary segmentation method and established its consistency. Their Assumption 3 on detection requirement translates to our setting as
@@
Ignoring logarithmic factors, our detection requirement (5) is weaker than (A.1) by a factor of $p/\min(\|\bm{\Sigma}\|\sigma_{*}^{-1}\mathfrak{s},\mathfrak{s}_{*})$, with $\mathfrak{s}_{*}=|\bm{\Sigma}\bm{\delta}|_{0}$, and (8) is weaker by a factor of $\sqrt{p}$. Thus, with OcScan (Section 4.1), we achieve an improvement by a factor of
@@
It is important to note that the requirement (A.1) is minimax optimal, up to $\log(n)$, for the covariance change point problem in general (Wang et al., 2021a, , Lemma 3.1). The improved detection performance of OcScan arises from leveraging the specific structure of the linear model, which also enables the careful analysis of the phase transition with respect to the inherent sparsity $\mathfrak{s}$. Additionally, the estimation rate achieved by our refined procedure described in Section 4.2, nearly matches the minimax lower bound of order $\kappa^{-2}B^{2}$ reported in Wang et al., 2021a (, Lemma 3.2).
Moen, (2024) investigated the minimax optimality in covariance change point testing over the class of parameters that are $d$-sparse for any $d\in[p]$, defined as
$\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0}):=\sup_{\begin{subarray}{c}\mathbf{v}:\,|\mathbf{v}|_{2}=1,\\ |\mathbf{v}|_{0}=d\end{subarray}}|\mathbf{v}^{\top}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})\mathbf{v}|,$
which implies that the leading eigenvector of $\bm{\Sigma}_{1}-\bm{\Sigma}_{0}$ is sparse with $d$ non-zero elements. Their (computationally intractable) test matches the minimax lower bound on the detection boundary, which requires
$\left(\frac{\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})}{\Lambda^{d}_{\max}(\bm{\Sigma}_{0})\vee\Lambda^{d}_{\max}(\bm{\Sigma}_{1})}\right)^{2}\Delta\;\gtrsim\;d\log(ep/d)\vee\log\log(8n)$ (A.3)
for some $d$. Some elementary calculations show that
$\Lambda^{d}_{\max}(\bm{\Sigma}_{0})\vee\Lambda^{d}_{\max}(\bm{\Sigma}_{1})\geq\max\left\{\Psi^{2},\max_{i\in[p]}(\bm{\Sigma})_{i,i}\right\},$
$\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})\leq\|\bm{\Sigma}_{1}-\bm{\Sigma}_{0}\|\leq\|\bm{\Sigma}_{1}-\bm{\Sigma}_{0}\|_{F}\lesssim|\bm{\Sigma}\bm{\delta}|_{2}+|\bm{\beta}_{1}^{\top}\bm{\Sigma}\bm{\beta}_{1}-\bm{\beta}_{0}^{\top}\bm{\Sigma}\bm{\beta}_{0}|,$
where $(\bm{\Sigma})_{i,i^{\prime}}$ denotes the element in the $i$th row and $i^{\prime}$th column of $\bm{\Sigma}$. To facilitate the comparison between (A.3) and our results, we relate the notion of $d$-sparsity and the sparsity of $\bm{\Sigma}\bm{\delta}$, as $d=\mathfrak{s}_{*}+1$. Then, again assuming (A.2), our requirement that either (5) or (8) holds, is weaker than (A.3) by a factor of $\max(d/\sqrt{p},1)$, save logarithmic factors. That is, the detection boundary of OcScan is more favourable in the scenario where $d\geq\sqrt{p}$.
Finally, Dette et al., (2022) investigated the covariance change point estimation problem by imposing the sparsity directly on $\bm{\Sigma}_{1}-\bm{\Sigma}_{0}$. For estimation consistency, their proposed method requires that $|\bm{\Sigma}_{j}|_{\infty}\lesssim 1$, $(\bm{\Sigma}_{j})_{i,i}\lesssim 1$, for $i\in[p+1]$ and $j\in\{0,1\}$, and
@@ B.1.1 Proof of Theorem 1
Proof of (i).
@@
Let $\mathbf{1}^{\pm}_{S}$ denote a $p$-dimensional vector with its $i$-th entry being 111 or 1-1- 1 if $i\in S$ and being zero if $i\not\in S$, and introduce $\mathcal{U}=\bigl{\{}\bm{u}=({\kappa}/{\sqrt{\mathfrak{s}}})\mathbf{1}^{\pm}_{S}\in\mathbb{R}^{p}:S\subseteq[p],\;|S|=\mathfrak{s}\bigr{\}}$, whose cardinality is $2^{\mathfrak{s}}{p\choose\mathfrak{s}}$. Define
$\mathsf{P}_{1}=2^{-\mathfrak{s}}{p\choose\mathfrak{s}}^{-1}\sum_{\bm{u}\in\mathcal{U}}\mathsf{P}_{\Delta,\bm{u},\bm{0}}\quad\text{and}\quad\mathsf{P}_{2}=2^{-\mathfrak{s}}{p\choose\mathfrak{s}}^{-1}\sum_{\bm{u}\in\mathcal{U}}\mathsf{P}_{n-\Delta,\bm{0},\bm{u}}.$
@@
Note that $\mathsf{P}_{\Delta,\bm{u},\bm{0}}$ and $\mathsf{P}_{n-\Delta,\bm{0},\bm{u}}$ lie in $\mathcal{P}_{\mathfrak{s},\sigma}^{\tau,n,p}$ under proper choices of $\kappa$ and $\Delta$, see (B.1) below. By the method of fuzzy hypotheses (see e.g. Section 2.7.4 in Tsybakov,, 2009), we have
$\inf_{\widehat{\theta}}\;\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{\mathfrak{s},n,p}(\tau)}\;\mathsf{P}\left(\bigl{|}\widehat{\theta}-\theta(\mathsf{P})\bigr{|}\geq\Delta\right)\;\geq\;\frac{1-\sqrt{2\alpha}}{2}\qquad\text{with}\quad\alpha:=\chi^{2}(\mathsf{P}_{1},\mathsf{P}_{0})=\chi^{2}(\mathsf{P}_{2},\mathsf{P}_{0}),$
where $\mathsf{P}_{0}$ denotes the distribution of $\{(\mathbf{x}_{t},Y_{t})\}_{t\in[n]}$ such that $\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\mathbf{I})$ and independently $Y_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2})$. Define $\mathsf{P}_{\bm{u}}$ as in Lemma B.2, and $Z_{k}$ as the sum of $k$ independent Rademacher random variables. Then, by Lemma B.2, we have
$\alpha+1$ $=2^{-2\mathfrak{s}}{p\choose\mathfrak{s}}^{-2}\sum_{\bm{u},\bm{v}\in\mathcal{U}}\left[\mathsf{E}_{\mathsf{P}_{\bm{0}}}\left(\frac{\mathrm{d}\mathsf{P}_{\bm{u}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\frac{\mathrm{d}\mathsf{P}_{\bm{v}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)\right]^{\Delta}$
=($={p\choose\mathfrak{s}}^{-2}\sum_{S,T\subseteq[p],\,|S|=|T|=\mathfrak{s}}\mathsf{E}\left[\left(\left(1-\frac{Z_{|S\cap T|}}{\mathfrak{s}}\frac{\kappa^{2}}{\sigma^{2}}\right)^{2}-\left(\frac{\kappa^{2}}{\sigma^{2}}\right)^{2}\right)^{-\Delta/2}\right]$
=($={p\choose\mathfrak{s}}^{-1}\sum_{S\subseteq[p],\,|S|=\mathfrak{s}}\mathsf{E}\left[\left(\left(1-\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\frac{\kappa^{2}}{\sigma^{2}}\right)^{2}-\left(\frac{\kappa^{2}}{\sigma^{2}}\right)^{2}\right)^{-\Delta/2}\right].$
Let $\rho:=\kappa^{2}/\sigma^{2}$ and consider $\rho\leq 1/4$. Then, by Lemma B.1,
$\mathsf{E}\left[\left(\left(1-\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\frac{\kappa^{2}}{\sigma^{2}}\right)^{2}-\left(\frac{\kappa^{2}}{\sigma^{2}}\right)^{2}\right)^{-\Delta/2}\right]$ $\leq\mathsf{E}\left[\left(1-2\rho\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}-\rho^{2}\right)^{-\Delta/2}\right]$
$\leq\exp(\Delta\rho^{2})\mathsf{E}\left[\left(1-\frac{2\rho}{1-\rho^{2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\right)^{-\Delta/2}\right].$
By symmetry of Rademacher distribution, we can find an event $A$ such that $\mathsf{P}(A)=1/2$, $A\subseteq\{Z_{|S\cap[\mathfrak{s}]|}\geq 0\}$ and $A^{c}\subseteq\{Z_{|S\cap[\mathfrak{s}]|}\leq 0\}$. As $2\rho/(1-\rho^{2})\leq 8/15\leq 3/4$, we apply again Lemma B.1 and obtain
$\mathsf{E}\left[\left(1-\frac{2\rho}{1-\rho^{2}}\cdot\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\right)^{-\Delta/2}\right]$
[…]
=\;= $\frac{\sqrt{2}}{2}\left[\cosh\left(\frac{2\Delta}{\mathfrak{s}}\cdot\frac{2\rho}{1-\rho^{2}}\right)\right]^{|S\cap[\mathfrak{s}]|/2}+\frac{\sqrt{2}}{2}\left[\cosh\left(\frac{\Delta}{\mathfrak{2s}}\cdot\frac{2\rho}{1-\rho^{2}}\right)\right]^{|S\cap[\mathfrak{s}]|/2}$
[…]
where the second inequality is by Hölder’s inequality. Note that the summation involving $|S\cap[\mathfrak{s}]|$ can be seen as the expectation of the hypergeometric distribution with parameters $(p,\mathfrak{s},\mathfrak{s})$. Let $\xi\sim\mathrm{Bin}(\mathfrak{s},\mathfrak{s}/p)$. Combining the above calculation and applying Lemma B.3, we obtain
$\alpha+1\leq\;$ $\sqrt{2}\exp(\Delta\rho^{2})\cdot\mathsf{E}\left[\cosh\left(\frac{2\Delta}{\mathfrak{s}}\cdot\frac{2\rho}{1-\rho^{2}}\right)^{\xi/2}\right]$
=\;= $\sqrt{2}\exp(\Delta\rho^{2})\cdot\exp\left[\mathfrak{s}\log\left(1+\frac{\mathfrak{s}}{p}\left\{\cosh\left(\frac{2\Delta}{\mathfrak{s}}\cdot\frac{2\rho}{1-\rho^{2}}\right)^{1/2}-1\right\}\right)\right]$
[…]
We set $\rho$ such that
$\frac{5\Delta\rho}{\mathfrak{s}}=\log\left(1+\frac{p}{\mathfrak{s}^{2}}\log\left(\frac{17}{12\sqrt{2}}\right)+\sqrt{\left(1+\frac{p}{\mathfrak{s}^{2}}\log\left(\frac{17}{12\sqrt{2}}\right)\right)^{2}-1}\right)$ (B.1)
It together with $\mathfrak{s}\log\bigl{(}1+\tfrac{p}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p}}{\mathfrak{s}}\bigr{)}\leq\sqrt{\tfrac{25}{2}\log\left(\tfrac{17}{12\sqrt{2}}\right){\Delta}}$ implies
$\Delta\rho^{2}=\frac{\Delta^{2}\rho^{2}}{\Delta}\leq\frac{\mathfrak{s}^{2}\log^{2}\left(1+\frac{p}{\mathfrak{s}^{2}}\vee\frac{\sqrt{p}}{\mathfrak{s}}\right)}{25\Delta}\leq\frac{1}{2}\log\left(\frac{17}{12\sqrt{2}}\right).$
Further, by the identity $\cosh\bigl{(}\log(1+x+\sqrt{x^{2}+2x})\bigr{)}=1+x$ and the inequality $\sqrt{1+x}\leq 1+x/2$, for $x\geq 0$, we have
$\alpha+1\leq\sqrt{2}\exp(\Delta\rho^{2})\cdot\left(1+\frac{1}{2\mathfrak{s}}\log\left(\frac{17}{12\sqrt{2}}\right)\right)^{\mathfrak{s}}\leq\frac{17}{12},$
which leads to $\alpha\leq 5/12$, and thus
@@
$\inf_{\widehat{\theta}}\;\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{\mathfrak{s},n,p}(\tau)}\;\mathsf{P}\left(\bigl{|}\widehat{\theta}-\theta(\mathsf{P})\bigr{|}\geq\Delta\right)\;\geq\;\frac{1-\sqrt{2\alpha}}{2}\;\geq\;\frac{1-\sqrt{5/6}}{2}\;>\;\frac{1}{25}.$
Since $\log(1+ux)\geq u\log(1+x)$ for $x\geq 0$ and $u\in[0,1]$, we derive from (B.1) that
$\frac{\Delta\rho}{\mathfrak{s}}\geq\frac{2}{5}\log\left(\frac{17}{12\sqrt{2}}\right)\log\left(1+\frac{p}{\mathfrak{s}^{2}}\vee\frac{\sqrt{p}}{\mathfrak{s}}\right).$
Proof of (ii).
We adopt the notation from Part (i) and consider now the following fuzzy hypotheses
$\mathsf{P}_{1}=2^{-\mathfrak{s}}{p\choose\mathfrak{s}}^{-1}\sum_{\bm{u}\in\mathcal{U}}\mathsf{P}_{\Delta,\bm{u},\bm{0}}\quad\text{and}\quad\mathsf{P}_{2}=2^{-\mathfrak{s}}{p\choose\mathfrak{s}}^{-1}\sum_{\bm{u}\in\mathcal{U}}\mathsf{P}_{\Delta+2\gamma,\bm{u},\bm{0}},$
where $\Delta\rho=\tau\mathfrak{s}\log\bigl{(}1+\tfrac{p}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p}}{\mathfrak{s}}\bigr{)}$ and $\gamma\leq\Delta/2$. Then, similarly, we have
$\inf_{\widehat{\theta}}\;\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{\mathfrak{s},n,p}(\tau)}\;\mathsf{P}\left(\bigl{|}\widehat{\theta}-\theta(\mathsf{P})\bigr{|}\geq\gamma\right)\;\geq\;\frac{1-\sqrt{\alpha/2}}{2},$
where $\alpha=\chi^{2}(\mathsf{P}_{2},\mathsf{P}_{1})$ and thus
@@
$\alpha\;=\;{p\choose\mathfrak{s}}^{-1}\sum_{S\subseteq[p],\,|S|=\mathfrak{s}}\left(\left(1-\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\rho\right)^{2}-\rho^{2}\right)^{-\gamma}-1.$
Again by the similar calculation as in Part (i), we have
$\alpha+1\;\leq\;\sqrt{2}\exp(2\gamma\rho^{2})\cdot\exp\left[\mathfrak{s}\log\left(1+\frac{\mathfrak{s}}{p}\left\{\cosh\left(\frac{10\gamma\rho}{\mathfrak{s}}\right)^{1/2}-1\right\}\right)\right].$
We set $\gamma$ such that

### v3

@@ Appendix A Connection to the covariance change detection problems
A.1 Change size measured in operator norm
Wang et al. (2021a) proposed a projection-based binary segmentation method and established its consistency. Their Assumption 3 on detection requirement translates to our setting as
@@ A.1 Change size measured in operator norm
Ignoring logarithmic factors, our detection requirement (5) is weaker than (A.1) by a factor of (at least) $p/\min(\|\bm{\Sigma}\|\sigma_{*}^{-1}\mathfrak{s},\mathfrak{s}_{*})$, with $\mathfrak{s}_{*}=|\bm{\Sigma}\bm{\delta}|_{0}$, and (8) is weaker by a factor of $\sqrt{p}$. Thus, with OcScan (Section 4.1), we achieve an improvement by a factor of
@@
It is important to note that the requirement (A.1) is minimax optimal, up to $\log(n)$, for the covariance change point problem in general (Wang et al., 2021a, Lemma 3.1). The improved detection performance of OcScan arises from leveraging the specific structure of the linear model, which also enables the careful analysis of the phase transition with respect to the inherent sparsity $\mathfrak{s}$. Additionally, the estimation rate achieved by our refinement method described in Section 4.2, nearly matches the minimax lower bound of order $\kappa^{-2}B^{2}$ reported in Wang et al. (2021a, Lemma 3.2).
A.2 Element-wise sparsity of the change
Dette et al. (2022) investigated the covariance change point estimation problem by imposing the sparsity directly on $\bm{\Sigma}_{1}-\bm{\Sigma}_{0}$. Let $(\bm{\Sigma})_{i,i^{\prime}}$ denotes the element in the $i$th row and $i^{\prime}$th column of $\bm{\Sigma}$. Then for estimation consistency, their proposed method requires that $|\bm{\Sigma}_{j}|_{\infty}\lesssim 1$, $(\bm{\Sigma}_{j})_{i,i}\lesssim 1$, for $i\in[p+1]$ and $j\in\{0,1\}$, and
@@ A.2 Element-wise sparsity of the change
A.3 Sparsity in the principal direction of the change
Moen (2024) investigated the minimax optimality in covariance change point testing over the class of parameters that are $d$-sparse for any $d\in[p]$, which fulfil
$\|\bm{\Sigma}_{1}-\bm{\Sigma}_{0}\|=\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0}):=\sup_{\begin{subarray}{c}\mathbf{v}:\,|\mathbf{v}|_{2}=1,\\ |\mathbf{v}|_{0}=d\end{subarray}}|\mathbf{v}^{\top}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})\mathbf{v}|,$
i.e. the leading eigenvector of $\bm{\Sigma}_{1}-\bm{\Sigma}_{0}$ is sparse with $d$ non-zero elements. Their (computationally intractable) test matches the minimax lower bound on the detection boundary, which requires
$\min\left\{\frac{\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})}{(\|\bm{\Sigma}_{0}\|\vee\|\bm{\Sigma}_{1}\|)-\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})},\left(\frac{\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})}{(\|\bm{\Sigma}_{0}\|\vee\|\bm{\Sigma}_{1}\|)-\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})}\right)^{2}\right\}\Delta$
$\gtrsim\;d\log(ep/d)\vee\log\log(8n)$
for some $d$ which, combined with that $\bm{\Sigma}_{j},\,j=0,1$, are positive definite, simplifies to
$\left(\frac{\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})}{\|\bm{\Sigma}_{0}\|\vee\|\bm{\Sigma}_{1}\|}\right)^{2}\Delta\;\gtrsim\;d\log(ep/d)\vee\log\log(8n).$ (A.3)
Some elementary calculations show that
$\|\bm{\Sigma}_{0}\|\vee\|\bm{\Sigma}_{1}\|$ $\geq\max(\Psi^{2},\|\bm{\Sigma}\|),\text{ \ and }$
$\Lambda^{d}_{\max}(\bm{\Sigma}_{1}-\bm{\Sigma}_{0})$ $\leq\|\bm{\Sigma}_{1}-\bm{\Sigma}_{0}\|\leq\|\bm{\Sigma}_{1}-\bm{\Sigma}_{0}\|_{F}\lesssim|\bm{\Sigma}\bm{\delta}|_{2}+|\bm{\beta}_{1}^{\top}\bm{\Sigma}\bm{\beta}_{1}-\bm{\beta}_{0}^{\top}\bm{\Sigma}\bm{\beta}_{0}|.$
To facilitate the comparison between (A.3) and our results, we relate the notion of $d$-sparsity and the sparsity of $\bm{\Sigma}\bm{\delta}$, as $d=\mathfrak{s}_{*}+1$. Then, again assuming (A.2), our requirement that either (5) or (8) holds, is weaker than (A.3) by a factor of $\max(d/\sqrt{p},1)$, save logarithmic factors. That is, the detection boundary of OcScan is more favourable in the scenario where $d\geq\sqrt{p}$.
A.4 Variance changes in responses
Under the AMOC model (1), we observe that
$Y_{t}\sim_{\mathrm{ind.}}\begin{cases}\mathcal{N}(0,\sigma_{0}^{2}),&1\leq t\leq\theta,\\ \mathcal{N}(0,\sigma_{1}^{2}),&\theta+1\leq t\leq n,\end{cases}\qquad\text{where}\qquad\sigma_{i}^{2}=\bm{\beta}_{i}^{\top}\bm{\Sigma}\bm{\beta}_{i}+\sigma^{2},$
which indicates that the change point $\theta$ is detectable as a change in $\mathsf{Var}(Y_{t})$. However, this strategy is not generally recommended, as it is completely powerless when $\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{0}\rvert_{2}=\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{1}\rvert_{2}$, even though $\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}\neq 0$, where $\bm{\delta}=\bm{\beta}_{1}-\bm{\beta}_{0}$.
Nevertheless, we relate our detection conditions in (5) and (8) to the minimax detection boundary for univariate variance change point detection (see Moen, 2024, Theorem 1),
$\Delta\Biggl\{\Bigl(\frac{\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert}{\min\{\sigma_{0}^{2},\sigma_{1}^{2}\}}\Bigr)\wedge\Bigl(\frac{\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert}{\min\{\sigma_{0}^{2},\sigma_{1}^{2}\}}\Bigr)^{2}\Biggr\}\asymp\log\log(n).$ (A.4)
For ease of comparison, assume that $\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{0}\rvert_{2}\neq\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{1}\rvert_{2}$ (or equivalently, $\sigma_{0}^{2}\neq\sigma_{1}^{2}$) and that $\bm{\Sigma}$ has its spectrum bounded away from zero and infinity. Under the latter condition, OcScan detects the change point $\theta$ whenever
$\Delta\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}^{2}\gtrsim\Psi^{2}\min\left\{\mathfrak{s}\log\bigl(p\log(n)\bigr),\,\sqrt{p\log\log(n)}\right\},$ (A.5)
where $\Psi=\max\{\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{0}\rvert_{2},\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{1}\rvert_{2},\sigma\}$ and $\mathfrak{s}=\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{0}$. Then, we distinguish the following two regimes.
  1. (a)
    Comparable variances, i.e. $\sigma_{0}^{2}\asymp\sigma_{1}^{2}$. In this case, (A.4) is equivalent to
    $\Delta\Bigl(\frac{\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert}{\sigma_{0}^{2}+\sigma_{1}^{2}}\Bigr)^{2}\asymp\log\log(n).$
    We relate $\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert$ to $\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}$ as follows:
    $\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert$ $=\Bigl\lvert\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{1}\rvert_{2}^{2}-\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{0}\rvert_{2}^{2}\Bigr\rvert$
    $=\lvert\bm{\delta}^{\top}\bm{\Sigma}(\bm{\beta}_{0}+\bm{\beta}_{1})\rvert$
    $\leq\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}\bigl(\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{0}\rvert_{2}+\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{1}\rvert_{2}\bigr)$ (A.6)
    $\leq 2\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}\,\Psi.$
    Since $\sigma_{0}^{2}+\sigma_{1}^{2}\asymp\Psi^{2}$, we obtain
    $\Delta\Bigl(\frac{\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert}{\sigma_{0}^{2}+\sigma_{1}^{2}}\Bigr)^{2}\lesssim\Delta\frac{\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}^{2}}{\Psi^{2}}.$
    Consequently, our requirement in (A.5) is always weaker than (A.4) when $p\lesssim\log\log(n)$, and is comparable when $\mathfrak{s}\asymp 1$ and $\log(p)\asymp\log\log(n)$. Otherwise, the boundary in (A.4) may be weaker, although this ultimately depends on the tightness of the inequality in (A.6).
  2. (b)
    Unbalanced variances, i.e. $\sigma_{0}^{2}\gg\sigma_{1}^{2}$ or $\sigma_{0}^{2}\ll\sigma_{1}^{2}$. WLOG, consider $\sigma_{0}^{2}\ll\sigma_{1}^{2}$. In this regime, (A.4) simplifies to
    $\Delta\frac{\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert}{\sigma_{0}^{2}}\asymp\log\log(n).$
    Because $\sigma_{0}^{2}\ll\sigma_{1}^{2}$ implies $\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert\asymp\sigma_{1}^{2}\asymp\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}^{2}$, we obtain
    $\Delta\frac{\lvert\sigma_{0}^{2}-\sigma_{1}^{2}\rvert}{\sigma_{0}^{2}}\asymp\Delta\frac{\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}^{2}}{\sigma_{0}^{2}}\gg\Delta\frac{\lvert\bm{\Sigma}^{1/2}\bm{\delta}\rvert_{2}^{2}}{\Psi^{2}}.$
    Hence, in this setting, the boundary (A.4) is weaker than (A.5) whenever $p\gtrsim\log\log(n)$.
Overall, this comparison suggests that additionally scanning for changes in $\mathsf{Var}(Y_{t})$ can be advantageous, particularly when $\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{0}\rvert_{2}$ and $\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{1}\rvert_{2}$ differ substantially. Note, however, that the change point literature for high-dimensional linear models (including the present paper) primarily focuses on the more challenging regime described in (a).
@@ B.1.1 Proof of Theorem 1
We define
$\mathcal{M}=\bigl\{\lfloor\log_{2}(n^{2\nu})\rfloor,\lfloor\log_{2}(n^{2\nu})\rfloor+1,\ldots,\lfloor\log_{2}(n/4)\rfloor\bigr\}$
and choose, with $\underline{c}_{1}=1/6$,
$\kappa^{2}=\underline{c}_{1}\sigma^{2}\mathfrak{s}\log\Bigl(1+\tfrac{p{\log\log(n)}}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p{\log\log(n)}}}{\mathfrak{s}}\Bigr).$
@@
Let $\mathbf{1}^{\pm}_{S}$ denote a $p$-dimensional vector with its $i$th entry being $1$ or 1-1 if $i\in S$ and being zero if $i\not\in S$, and introduce $\mathcal{U}=\bigl\{\bm{u}=({\kappa}/{\sqrt{\mathfrak{s}}})\mathbf{1}^{\pm}_{S}\in\mathbb{R}^{p}:S\subseteq[p],\;|S|=\mathfrak{s}\bigr\}$. Define
$\mathsf{P}^{(1)}=\frac{1}{|\mathcal{U}|\,|\mathcal{M}|}\sum_{\bm{u}\in\mathcal{U}}\sum_{m\in\mathcal{M}}\mathsf{P}_{2^{m},2^{-{m}/{2}}\bm{u},\bm{0}}\,.$
@@
Since $\mathsf{P}_{2^{m},2^{-{m}/{2}}\bm{u},\bm{0}}$ lies in $\mathcal{P}^{\mathfrak{s},n,p}_{\mathbf{I},\sigma}(\tau)$, by the method of fuzzy hypotheses (see e.g. Section 2.7.4 in Tsybakov, 2009), we have
@@
$\inf_{\psi}\biggl\{\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{0,n,p}}\mathsf{E}_{\mathsf{P}}\left[\psi\left((Y_{t},\mathbf{x}_{t})_{t\in[n]}\right)\right]+\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{\mathfrak{s},n,p}(\tau)}\mathsf{E}_{\mathsf{P}}\left[1-\psi\left((Y_{t},\mathbf{x}_{t})_{t\in[n]}\right)\right]\biggr\}\geq\frac{\exp\left(-\chi^{2}(\mathsf{P}^{(1)},\mathsf{P}^{(0)})\right)}{4},$
where $\mathsf{P}^{(0)}\in\mathcal{P}_{\mathbf{I},\sigma}^{0,n,p}$ is the distribution of $\{(\mathbf{x}_{t},Y_{t})\}_{t\in[n]}$ such that $\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\mathbf{I})$ and independently $Y_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2})$. Define $\mathsf{P}_{\bm{u}}$ as in Lemma B.2, and $Z_{k}$ as the sum of $k$ independent Rademacher random variables. Then, by Lemma B.2, we have
$\chi^{2}(\mathsf{P}^{(1)},\mathsf{P}^{(0)})+1$
$=\,$ $\frac{1}{|\mathcal{U}|^{2}\,|\mathcal{M}|^{2}}\sum_{m,\ell\in\mathcal{M}}\sum_{\bm{u},\bm{v}\in\mathcal{U}}\left[\mathsf{E}_{\mathsf{P}_{\bm{0}}}\left(\frac{\mathrm{d}\mathsf{P}_{2^{-{m}/{2}}\bm{u}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\frac{\mathrm{d}\mathsf{P}_{2^{-{\ell}/{2}}\bm{v}}}{\mathrm{d}\mathsf{P}_{\bm{0}}}\right)\right]^{2^{m}\wedge 2^{\ell}}$
@@
$=\,$ $\frac{1}{|\mathcal{M}|^{2}{p\choose\mathfrak{s}}}\sum_{m,\ell\in\mathcal{M}}\sum_{S\subseteq[p],\,|S|=\mathfrak{s}}\mathsf{E}\left[\left(\left(1-\frac{Z_{|S\cap[\mathfrak{s}]|}}{\mathfrak{s}}\frac{\kappa^{2}}{\sigma^{2}2^{(m+\ell)/2}}\right)^{2}-\left(\frac{\kappa^{2}}{\sigma^{2}2^{(m+\ell)/2}}\right)^{2}\right)^{-2^{(m\wedge\ell)-1}}\right].$
Let $\rho:=\kappa^{2}/\sigma^{2}$ and then $\rho\leq 2^{\lfloor\log_{2}(n^{2\nu})\rfloor}/3$. Then, by Lemma B.1, for $m,\ell\in\mathcal{M}$,